OpenStudy (anonymous):

The hypotenuse if a right triangle is 3cm linger than twice the length of the shorter leg. The longer leg measures 12cm . Find the lengths of the shorter leg and the hypotenuse

OpenStudy (anonymous):



OpenStudy (anonymous):

I can't see the while pic

OpenStudy (anonymous):

Whole

OpenStudy (anonymous):

what part can't you see?

OpenStudy (anonymous):

Half, i only see 12 abd letter s

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

oh okay. the hypotenuse has 3+2s on it. S is your base, and 12 is the height/larger leg

OpenStudy (anonymous):

What will be the equation? 3+2s+s=12?

OpenStudy (anonymous):

Not exactly. You would have to be using the Pythagorean theorem I think. so A^2 + B^2 = C^2. A^2 is 144, B^2 is the same as saying S^2, and C^2 = 4s^2 + 12s + 9

OpenStudy (anonymous):

So entire equation put together: 144 + S^2 = 4s^2 + 12s + 9

OpenStudy (anonymous):

Oh wait i'll try to solve

OpenStudy (anonymous):

I got x=9?

OpenStudy (anonymous):

That is what I got

OpenStudy (anonymous):

So the hypotenuse is 21 and shorter leg is 9?

OpenStudy (anonymous):

yes